### v1

@@ 3.1.2 Settings
To evaluate the effectiveness of the proposed method, several experiments have been conducted. We divide the dataset into training, validation, and testing sets, resulting in 3,686 cells for training, 922 cells for validation, and 1,153 cells for testing. For all experiments, we preprocess the images by resizing them to a size of $3\times 224\times 224$ to align with the ResNet-18 input shape, followed by normalization; and preprocess the tabular data by applying standard scaler from the Scikit-learn library. To investigate the impact of feature extraction, we perform the proposed method on both Protocol 1 and Protocol 2.
@@ 3.2 Training Methodology
In this study, an Adam optimizer is used with a learning rate of 0.0005 to optimize the MSE objective function (Section 2.2). The models are trained with a batch size of 40 in 100 epochs using the Pytorch framework. The network with the highest Spearman correlation on the validation dataset is selected and continuously tested on the testing dataset. All models are trained on a GeForce RTX 3090 GPU.
3.3 Evaluation Metrics
The model performance is evaluated and measured using four metrics, including Spearman correlation, MAE, MSE, and $R^{2}$ score. First, Spearman correlation [13] is a non-parametric statistical method to assess the relationships between two variables based on ranks. To calculate this coefficient, we first order the two variables $\hat{y}$ and $y$ from least to greatest. Let $D^{(i)}$ be the distance between the two corresponding ranks. The Spearman correlation is computed using the formula
$r_{S}=1-\frac{6{\sum_{i=1}^{N}{D^{(i)}}^{2}}}{N(N^{2}-1)},$ (2)
where $i=1,...,N$.
Secondly, MAE is the mean of the difference between $\hat{y}$ and $y$ in absolute value, calculated as
$MAE=\frac{1}{N}{\sum_{i=1}^{N}|y^{(i)}-\hat{y}^{(i)}|}.$ (3)
MSE is the mean of the square of the difference between $\hat{y}$ and $y$ as the equation
$MSE=\frac{1}{N}{\sum_{i=1}^{N}|y^{(i)}-\hat{y}^{(i)}|^{2}}.$ (4)
$R^{2}$ (also called the coefficient of determination) [23] is defined as the fraction of the variation in the dependent variable $\hat{y}$ that is predicted from the independent variable $y$. It has the worst value of $\infty$- ∞, the best value of $+1$ [5], and can be calculated with the formula
$R^{2}=1-\frac{\sum_{i=1}^{N}|y^{(i)}-\hat{y}^{(i)}|^{2}}{\sum_{i=1}^{N}|y^{(i)}-\overline{y}|^{2}},$ (5)
where $\overline{y}$ is the mean value of $y$.
4 Result & Discussion
4.1 Model Performance
Table 3: Result summary of four approaches on the testing dataset
Protocol
Approach Model
@@ 4.1 Model Performance
1 Combined model 0.689 0.373 0.249 0.483
1 2 SarcNet 0.825 0.310 0.159 0.671
3 Combined model 0.809 0.320 0.180 0.627
2 4 SarcNet 0.831 0.310 0.161 0.668
In this section, we compare the performance of SarcNet with a combined model, which integrates the strengths of both the ResNet-18 module and linear regression to enhance the prediction performance of the sarcomere structure organization score. Given an input image of a single cell, the ResNet-18 module directly processes the raw image, while linear regression utilizes a set of extracted features. Both component models generate a predicted score ranging from one to five. The final score is computed as the mean of two outputs derived from the individual predictions from the ResNet-18 module and linear regression.
To understand the effectiveness of feature extraction and deep-learning models on quantifying sarcomere structure organization, Table 3 provides the four performance metrics, including Spearman correlation, MAE, MSE, and $R^{2}$ score, of four approaches on the testing dataset. The last two approaches (3, 4), which use Protocol 2, claim much superior outcomes (0.809-0.831) compared to those with Protocol 1 (0.689-0.825). A highlighted point is that while the combined model with Protocol 2 dramatically outperforms the performance in Protocol 1, the SarcNet model reveals nearly identical results in all performance metrics (Spearman: 0.83, MAE: 0.31, MSE: 0.16, and $R^{2}$ score: 0.67) between the two different methods. An explanation for this could be that the linear regression model with fewer features leads to poor performance, which significantly affects the performance of the combined model, lowering the results from 0.809 to 0.689.
In addition, the SarcNet model claims strength in extracting useful information from input images and features compared to the combined model. While the combined model requires the six features extracted from the deep-learning model to leverage the performance (Table 3), the former can forgo this step and still maintain similar outcomes. This would effectively speed up the framework and minimize the memory storage in assessing sarcomere structure organization in hiPSC-CMs.
Figure 3: Visualization results for the GT and the corresponding predictions from the SarcNet model. The first row illustrates the predicted examples with $MAE\leq 0.1$, the second row with $0.1<MAE\leq 1$, and the last row with $MAE>1$.
Fig. 3 presents prediction examples with different ranges of MAE. It is noticeable that the predictions from the proposed architecture significantly reach the original targets, especially those from the first row. The last row, however, shows some incorrectly predicted results. In particular, Fig. 3(g) is mistakenly predicted since this cell has a section with considerably organized z-discs, which confuses the model. Furthermore, Fig. 3(h) is enlarged too much, making the alpha-actinin-2 patterns in this image appear crisper with higher density, resulting in a higher prediction score. In contrast, Fig. 3(i) is shrunk too tiny relative to the original, resulting in blurrier alpha-actinin-2 patterns and a substantially lower prediction score.
4.2 Model Interpretation
Figure 4: Visualization of Grad-CAM heatmaps for different examples.
@@ 4.2 Model Interpretation
Grad-CAM heatmaps (Fig. 4) are used to interpret the SarcNet model predictions. Fig. 4(c), (d), and (e) produce the highest values focusing mainly on the sarcomere structure, while Fig. 4(b) has the lowest value at the center region of the cell. An explanation for this could be that the two sides of the cell have clearer fiber patterns while the center does not; as a result, it would be easier for the model to make decisions based on the two sides of the cell. In contrast, Fig. 4(a) illustrates a mistaken case in which the model seems to focus more on the regions with organized z-discs than the diffuse regions.
4.3 Changes in Cell Organization Over Time
Figure 5: Histogram of the predicted scores for day 18 (blue) and day 32 (red).
@@ 4.3 Changes in Cell Organization Over Time
Fig. 5 aims to confirm the pattern of increasing organization levels between days 18 and 32. It is noticeable that the predicted scores from the SarcNet model are consistent with the culture period and expert annotations, with a slight shift to the right of the 32-day histogram. While most of the hiPSC-CM organization on day 18 concentrates around level 3, with most of the regions being disorganized puncta, day 32 indicates more organized regions with level 4. However, most of the cells on day 32 are still immature and need further culture. This finding agrees with the research in [6], which has investigated the culture of hiPSC-CMs on days 30, 90, and 200 and proclaimed that cell contractility reaches the maximum level at the 200-day time point.
4.4 Comparison
Table 4: Comparison of other models on the testing dataset
Model
Feature
extraction
Spearman
Correlation
MAE MSE
$R^{2}$
score
ResNet-18 [11] No 0.829 0.453 0.300 0.379
DenseNet [14] No 0.813 0.340 0.194 0.599
AlexNet [16] No 0.788 0.346 0.195 0.596
Linear Regression [7] Protocol 2 0.756 0.373 0.234 0.514
Ours Protocol 2 0.831 0.310 0.161 0.668
Table 4 compares the performance of our proposed framework to other models and claims a noticeable result. As can be seen from the table, the ResNet-18 model and SarcNet achieve an approximately similar Spearman correlation of 0.829 and 0.831, respectively, but are significantly different in the results of the other metrics. An explanation for this finding could be that the Spearman correlation only measures the association based on ranks between the GT and predictions but does not consider the distance between them as MAE and MSE. As a result, the ResNet-18 model seems to quantify the sarcomere structure organization in the correct rank of scores but still achieves poor performance in MAE and MSE (0.453 and 0.300, respectively). In contrast, thanks to feature extraction, SarcNet can get more relevant information and score the sarcomere structure with smaller values in MAE and MSE, which are 0.310 and 0.161, respectively.
4.5 Limitations and Future Directions
@@
Some limitations in this work need addressing before quantifying sarcomere structure organization. One of the most noteworthy is the lack of single-cell segmentation progress. Experts manually mark single-cell boundaries in the AISC dataset [7], as there is no accessible automated framework. This work should be considered challenging because the individual cells are not completely separated from one another but overlap; also, high cell density regions should be avoided. In the future, we intend to develop an automated procedure that addresses this challenge and facilitates more accurate quantification of sarcomere structure organization.
5 Conclusion
This work highlights key discoveries in single-cell imaging of hiPSC-CMs. First, in this study, we develop SarcNet, a deep learning-based method for quantifying sarcomere structure organization. In particular, we propose this model to leverage predictions by concatenating the ResNet-18 module output with a representation vector of quantitative single-cell measurements of subcellular organization. Second, we conduct extensive experiments and compare SarcNet to other approaches to illustrate the usefulness of the proposed method, with the best performance in Spearman correlation of 0.831. Finally, we speed up the training and interference processes by reducing the number of features while retaining a performance of 0.825.
Conflict of Interest
The authors have no conflicts of interest to declare.
Acknowledgment
This study is supported by the VinUni-Illinois Smart Health Center, VinUniversity.

### v2

@@ 3.1.2 Experimental Settings
To evaluate the effectiveness of the proposed method, several experiments have been conducted. We divide the dataset into training, validation, and testing sets with 3,686 cells for training, 922 cells for validation, and 1,153 cells for testing. For all experiments, we preprocess the images by resizing them to a size of $3\times 224\times 224$ to align with the ResNet-18 input shape, followed by normalization; and preprocess the tabular data by applying standard scaler from the Scikit-learn library. To investigate the impact of feature extraction, we perform the proposed method on both Protocol 1 and Protocol 2.
@@ 3.2 Training Methodology
In this study, an Adam optimizer was used with a learning rate of 0.0005 to optimize the MSE objective function (Section 2.1). The model was trained with a batch size of 40 in 100 epochs using the Pytorch framework. The network with the highest Spearman correlation on the validation dataset was selected and continuously tested on the testing dataset. All models were trained on a GeForce RTX 3090 GPU.
3.3 Experimental Results
To quantify the effectiveness of our model, we compares the performance of SarcNet with linear regression used in [10] and other state-of-the-art CNN models on the test set. Table 1 shows that SarcNet using both feature extraction protocols outperform other models, with a 7.5%, 0.063, 0.075, and 15.7% improvement compared to linear regression in Spearman correlation, MAE, MSE, and $R^{2}$ score, respectively. It is noticeable that the ResNet-18 model and SarcNet achieve an approximately similar Spearman correlation of 0.829 and 0.831, but are significantly different in the results of the other metrics. An explanation for this finding could be that the Spearman correlation only measures the association based on ranks between the GT and predictions but does not consider the distance between them as MAE and MSE. The ResNet-18 model seems to quantify the sarcomere structure organization in the correct rank of scores but still achieves poor performance in MAE and MSE. In contrast, thanks to feature extraction, SarcNet can get more relevant information and score the sarcomere structure with smaller values in MAE and MSE. The visualizations of predicted results by SarcNet are shown in Fig. 2A. Notably, in both feature extraction protocols, the SarcNet model reveals nearly identical results. This results indicate that the Protocol 1, which is more simple, also contributes significantly to the overall performance. The training and inference processes therefore can be speed up by simplifying feature extraction protocol.
Table 1: Quantitative assessment of SarcNet and other models on the testing dataset.
Model
@@ 3.3 Experimental Results
Linear Regression 0.756 0.373 0.234 0.514
ResNet-18 0.829 0.453 0.300 0.379
DenseNet 0.813 0.340 0.194 0.599
AlexNet 0.788 0.346 0.195 0.596
SarcNet (Protocol 1) 0.825 0.310 0.159 0.671
SarcNet (Protocol 2) 0.831 0.310 0.161 0.668
4 Discussion and Conclusion
@@
Grad-CAM heatmaps (Fig. 2B) are used to interpret the SarcNet model predictions. Fig. 2B(c), (d), and (e) produce the highest values focusing mainly on the sarcomere structure, while Fig. 2B(b) has the lowest value at the center region of the cell. An explanation for this could be that the two sides of the cell have clearer fiber patterns while the center does not; as a result, it would be easier for the model to make decisions based on the two sides of the cell. However, Fig. 2B(a) illustrates a mistaken case in which the model seems to focus more on the regions with organized z-discs than diffuse regions.
Fig. 2: Visualizations of examples of prediction result (A) and Grad-CAM heatmaps (B).
@@
Regarding changes in cell organization over time, Fig. 3 aims to confirm the pattern of increasing organization levels between days 18 and 32. It is noticeable that the predicted scores from the SarcNet model are consistent with the culture period, with a slight shift to the right of the 32-day histogram. While most of the hiPSC-CM organization on day 18 concentrates around level 3, with most of the regions being disorganized puncta, day 32 indicates more organized regions with level 4. However, most of the cells on day 32 are still immature and need further culture. This finding agrees with the research in [15], which proclaimed that cell contractility reaches the maximum level at the 200-day time point.
Fig. 3: Histogram of the predicted scores for day 18 (blue) and day 32 (red).
@@
To conclude, we developed SarcNet, a deep learning-based method for quantifying sarcomere structure organization. Experiments show that SarcNet outperforms other approaches, with the best performance in Spearman correlation of 0.831. It is also noticeable that we can reduce the number of features while retaining the performance. Some limitations need addressing before quantifying sarcomere structure organization at single cell level, with one of the most noteworthy is the lack of single-cell segmentation progress. Experts manually mark single-cell boundaries in the AISC dataset [10], as there is no accessible automated framework. This work is challenging because the individual cells are not completely separated from one another but overlap. Future work aims to develop an automated procedure that addresses this challenge and facilitates more accurate quantification of sarcomere structure organization.
5 Compliance with ethical standards
This research study was conducted retrospectively using human subject data made available in open access by [10]. Ethical approval was not required as confirmed by the license attached with the open access data.